OpenStudy (anonymous):

Intergral (2(x^4)^1/2-x^-1/2)dx= ?

OpenStudy (anonymous):

I have no idea what the actual function looks like based on what you posted...

OpenStudy (anonymous):

[2(x^4)^(1/2)-x^(-1/2]*dx= ?

OpenStudy (anonymous):

So rewrite it to something that makes it easier or more obvious for that matter to integrate.

OpenStudy (kropot72):

\[\int\limits_{}^{}(2*(x ^{4})^{-1/2}-x ^{-1/2})dx\] Is this it?

OpenStudy (anonymous):

I think the first term has a positive 1/2 exponent as in x^4 is in a square root. Whether the two is as well, I don't know.

OpenStudy (anonymous):

asker, why don't you just integrate it separately? i think what @kropot72 has is what you'ver written.

OpenStudy (kropot72):

If I give a positive sign to the 1/2 of the first term the result of integration is:\[2/3*x ^{2}-2\sqrt{x}+C\]

OpenStudy (kropot72):

If I give a negative sign to the 1/2 index of the first term the result of integration is:\[-(2/x)-2\sqrt{x}+C\]